# First-time assessment of glitch-induced bias and uncertainty in inference of extreme mass ratio inspirals

### v1

@@ III.1 Waveforms and glitches
Figure 1 shows the resulting SNR distribution for LISA-injected glitches from the LPF catalog: some $50\%$ have up to moderate amplitudes ($\rho\!\lesssim\!10$), while a minority exceed $\rho\gtrsim 10^{3}$. Very loud glitches such as these would likely be mitigated or masked in practice. In the next section, we use the annotated percentiles in Figure 1 to construct “mitigated” glitch backgrounds to be used in our estimation of inference bias and covariance.
@@ V Discussions and Conclusions
It is important to acknowledge the simplifying assumptions underlying this study. The FM formalism provides a computationally efficient approximation valid only in the high-SNR, near-linear regime, but may underestimate correlations or non-Gaussianity in the true posterior. Realistic LISA observations will also contain quieter EMRIs, additional and overlapping sources, and time-dependent instrumental noise. We note that LISA could experience other kinds of glitches such as sine-Gaussians whose morphology and statistics could further impact the parameter biases and covariances beyond what is reported here. This represents a potential avenue for future work.
@@ Appendix B Fisher matrix validation
Figures 4, 6, and 5 visually demonstrate that despite the ill-conditioned nature of EMRI FMs [Vallisneri_2008], the FM-derived posterior distributions closely match those obtained from MCMC sampling. This agreement indicates that our estimated FMs provide accurate local approximations to the likelihood, ensuring that the glitch-induced parameter fluctuations computed using Eq. 24 are not compromised by FM inaccuracies.

### v2

@@ III.1 Waveforms and glitches
In Figure 1, we visualize how glitches obscure EMRIs in the time-frequency domain. The diagonal tracks reveal the time and frequency evolution of the individual EMRI harmonics, while the sharp vertical lines indicate glitch placement over the 2 year long observation.
Figure 2 shows the SNR distribution of LISA-injected glitches from the LPF catalog: some $50\%$ have up to moderate amplitudes ($\rho\!\lesssim\!10$), while a minority exceed $\rho\gtrsim 10^{3}$. Very loud glitches such as these would likely be mitigated or masked in practice. In the next section, we use the annotated percentiles in Figure 2 to construct “mitigated” glitch backgrounds to be used in our estimation of inference bias and covariance.
Figure 1: A Short-Time Fourier Transform spectrogram of the TDI response of the prograde EMRI buried in an unmitigated glitch background (left), alongside the pure EMRI (right). We used a Tukey window of length $32,768$ samples and an overlap of $75\%$. This was to ensure a relatively high frequency resolution, making the EMRI modes (seen as arching diagonal lines) easy to distinguish without making the glitches (seen as sharp vertical lines) poorly resolved in time.
@@ V Discussions and Conclusions
It is important to acknowledge the simplifying assumptions underlying this study. The FM formalism provides a computationally efficient approximation valid only in the high-SNR, near-linear regime, but may underestimate correlations or non-Gaussianity in the true posterior. Realistic LISA observations will also contain quieter EMRIs, additional and overlapping sources, and time-dependent instrumental noise. We note that LISA could experience other kinds of glitches such as sine-Gaussians whose morphology and statistics could further impact the parameter biases and covariances beyond what is reported here.
We also reiterate that our study did not account for the $T$-channel contribution. Given the channel’s insensitivity to EMRIs but not glitches, we conjecture that including the $T$-channel would marginally increase the estimated biases. One other consideration is our assumption of exact glitch mitigation. In practice, glitch mitigation algorithms will leave non-zero residuals in the data which would further shift the biases. Investigating the impact of these residual-induced biases would be a fitting avenue for future research.
@@ Appendix B Fisher matrix validation
Figures 5, 7, and 6 visually demonstrate that despite the ill-conditioned nature of EMRI FMs [Vallisneri_2008], the FM-derived posterior distributions closely match those obtained from MCMC sampling. This agreement indicates that our estimated FMs provide accurate local approximations to the likelihood, ensuring that the glitch-induced parameter fluctuations computed using Eq. 24 are not compromised by FM inaccuracies.
@@
As a visual verification, Figure 8 shows matching biased posteriors for the prograde EMRI obscured by a particular glitch background mitigated to SNR 400.
@@
Figure 8: A comparison of the MCMC-derived and FM-derived posteriors for the prograde EMRI corrupted by a glitch background mitigated to SNR 400. We did not include stationary noise in the data.